A rectangular surface of side 10 cm and 15 cm is placed inside a uniform electric field of 25v/m such that the surface makes an angle 30 with the direction of electric field . Find the flux of the electric field through the rectangular surface

Grade:12

2 Answers

ritik
6 years ago
phi=E.A
SO PHI=E.A.COStheta
PHI=25.150.10-4.1.732/2=
CALCULATE IT AND GET YOUR ANSWER PLEASE APPROVE IF YOU GOT IT CLEARLY
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

E = 25V/m
A = 10×10^−2×15×10^−2
=1.5×10^−2m2
Now, we know that∫ϕ=∫Eda
=E.A
=Eacosθ
As the surface make300 but we have to calculate according to area vector which is perpendicular to the surface,∴→θ=900−300=600
∴ϕ = Eacosθ
= 25×1.5×10^−2×cos600
= 0.1875Nm2/C

Thanks and Regards